# Incomplete U-Statistics of Equireplicate Designs: Berry-Esseen Bound and Efficient Construction

### v1

@@ Abstract
U-statistics are a fundamental class of estimators that generalize the sample mean and underpin much of nonparametric statistics. Although extensively studied in both statistics and probability, key challenges remain: their high computational cost—addressed partly through incomplete U-statistics—and their non-standard asymptotic behavior in the degenerate case, which typically requires resampling methods for hypothesis testing. This paper presents a novel perspective on U-statistics, grounded in hypergraph theory and combinatorial designs. Our approach bypasses the traditional Hoeffding decomposition, the main analytical tool in this literature but one highly sensitive to degeneracy. By characterizing the dependence structure of a U-statistic, we derive a Berry–Esseen bound that applies to all incomplete U-statistics of deterministic designs, yielding conditions under which Gaussian limiting distributions can be established even in the degenerate case and when the order diverges. We also introduce efficient algorithms to construct incomplete U-statistics of equireplicate designs, a subclass of deterministic designs that, in certain cases, achieve minimum variance. Finally, we apply our framework to kernel-based tests that use Maximum Mean Discrepancy (MMD) and Hilbert–Schmidt Independence Criterion. In a real data example with CIFAR-10, our permutation-free MMD test delivers substantial computational gains while retaining power and type I error control.
@@ 1 Introduction
On the applications side, U-statistics have found widespread use across various fields. These range from classical statistical estimation and inference tasks: e.g., in nonparametric hypothesis testing (Bergsma and Dassios,, 2014; Yao et al.,, 2018), robust statistics (Joly and Lugosi,, 2016; Minsker and Wei,, 2020), bootstrap and resampling methods (Leucht and Neumann,, 2013; Bastian et al.,, 2024), to modern machine learning applications: e.g., in empirical risk minimization (Clémençon et al.,, 2008; Chen et al.,, 2023), supervised learning ensembles (Mentch and Hooker,, 2016; Peng et al.,, 2022) and kernel methods (Gretton et al.,, 2005, 2012; Liu et al.,, 2016). Second-order U-statistics are the most common and widely used in practice. This class includes estimators of variance, covariance, Kendall’s Tau, Spearman’s Rho, Wilcoxon statistics, Cramer-Von Mises statistics. Additionally, they encompass novel estimators of the Maximum Mean Discrepancy (MMD) and the Kernel Stein Discrepancy (KSD), which are kernel methods designed for two-sample testing (Gretton et al.,, 2012) and goodness-of-fit testing (Liu et al.,, 2016), respectively (see Schrab et al., (2022) for a discussion). Higher-order U-statistics are also relevant in applications, as they include estimators of dependence measures such as distance covariance (Székely et al.,, 2007), sign covariance (Bergsma and Dassios,, 2014), and the Hilbert–Schmidt Independence Criterion (HSIC) (Gretton et al.,, 2005).
@@
    To characterize the dependence structure of an incomplete U-statistic, we show that the dependency graph (Baldi and Rinott,, 1989) is the line graph of the design hypergraph.
@@
Organization. Section 2 reviews background material and sets the notation for the paper. Section 3 derives a Berry–Esseen bound for incomplete U-statistics of deterministic designs and establishes conditions for Gaussian limiting distributions, even in the degenerate case and when the order diverges. Section 4 presents efficient algorithms for constructing equireplicate designs. Section 5 validates our theoretical results through extensive simulations on kernel-based test statistics and a real-data application showcasing a permutation-free variant of the MMD test. Section 6 concludes with a summary and directions for future research. All proofs and technical details corresponding to each section of the main paper are provided in the supplementary material, where sections are labeled with an “S”. The R code for this paper can be found at https://github.com/CaesarXVII/IUstat_equireplicate_designs.
@@ 2.2 Equireplicate designs
When $k=2$, incomplete U-statistics based on equireplicate designs achieve minimum variance, as demonstrated by Theorem 1 on page 195 of Lee, (1990). When $k>2$, additional conditions are needed to ensue minimum variance, such as balanced incomplete block designs (BIBDs) and cyclic designs (see Lee, (1982), Examples 2 and 7, respectively). Further details are provided in S8.1.
The existence of equireplicate designs requires specific divisibility conditions. In general, given a nonzero integer $a$ and an integer $b$, we write $a\mid b$ to indicate that $a$ divides $b$, and $a\nmid b$ otherwise. For our design constructions, we use modular arithmetic and some basic group theory (see Gallian, (2021), ch. 2 for an overview on groups and modular arithmetic).
We write $b\pmod{n}$ to denote the remainder upon dividing $b$ by $n$, but set $b\pmod{n}\vcentcolon=n$ if the remainder is zero; we also write $\text{gcd}(b,n)$ as the greatest common divisor of $b$ and $n$. In Section 4, we present efficient algorithmic constructions of equireplicate designs that use these algebraic concepts. Since we relabeled 0 as $n$ when calculating values $\pmod{n}$, we consider $\mathbb{Z}_{n}=\{1,2,\ldots,n\}$ as the group of integers modulo $n$ with additive identity element $n$. This notation was chosen to ensure that the elements of our designs are labeled from the index set $V=\{1,\ldots,n\}$.
@@ 3.1 The dependency graph of incomplete U-statistics
In a dependency graph, there is an edge between two random variables only if they are dependent (absence of an edge implies independence). In the case of U-statistics, we include an edge between two random variables $h(S_{1})$ and $h(S_{2})$ if and only if $S\cap S_{2}\neq\emptyset$.
@@
In the following section, we show that a Gaussian limiting distribution can be established when there is a bound on the degree of the dependency graph, $\Delta(L(\mathcal{D}))$.
@@ 3.3 Asymptotic results in the finite and infinite order regimes
Moreover, from the Berry–Esseen bound in (8), it follows that under the assumptions of Theorem 2 with $p=3$, the convergence rate to the standard normal distribution is $n^{-1/2}$, up to logarithmic factors (see the proof of Theorem 2 for details). This rate matches the classical Berry–Esseen bound for the complete U-statistic in the finite-order non-degenerate case (see Lee, (1990), ch. 3.3 and references therein), again up to logarithmic terms. It is also only slightly slower than the $o(n^{-1/2})$ rate established in Korolyuk and Borovskikh, (1989) for the complete U-statistic in the finite-order degenerate case, where the limiting distribution is non-Gaussian. In that setting, the variance scales as $n^{k/2}$ for the complete statistic, while for the incomplete one it scales as $|D_{n}|^{1/2}$ (see Corollary 2). This last result confirms—and extends to deterministic designs—the findings of Remark 3.2 in Chen and Kato, (2019), which were stated for random designs.
@@
To apply Theorem 2 in the infinite-order regime, we need a consistent estimator of the variance when $k$ diverges. Existing works have proposed variance estimation methods for incomplete U-statistics in the infinite-order regime, which can be applied (Wang and Lindsay,, 2014; Song et al.,, 2019; Xu et al.,, 2024). However, specific details for implementation of our framework in the infinite-order regime is left for future work.
@@ 4 Efficient Construction of Equireplicate Designs
In this section, we highlight another major advantage of equireplicate designs, in addition to their asymptotic properties, which is that they can be constructed in linear time with respect to the design size.
@@
In the case of second-order incomplete U-statistics, we show that optimal equireplicate designs of any possible size can be efficiently constructed and have minimum variance. Therefore, equireplicate designs should be preferred when $k=2$, the most common and widely used case in practice.
When $k>2$, constructing $r$-equireplicate designs becomes challenging, as it relates to open problems in discrete mathematics. In Section 4.2 we present a construction which has still a linear computational complexity in the design size—when $k$ is fixed—and that allows $|D|=O(n^{2})$ (see Remarks S8 and 7 for further details). This construction, based on cyclic permutations (see Lee, (1982), Example 7), may be of independent interest for the combinatorial design and hypergraph community.
4.1 Construction of r-equireplicate designs when k = 2
@@ 5.1 MMD experiments
We observe that for $r\in\{\log(n),\,\log^{2}(n)\}$, the KS CIs lie below $Q_{0.975}$ even at $n=100$ and quickly contract toward $Q_{0.5}$ as $n$ increases. When $r=\log^{3}(n)$, a larger sample size is required before the upper bound drops below $Q_{0.975}$, but the decreasing trend is evident. In contrast, for larger values of $r$ (e.g., $r=n/2$), the distance to normality remains roughly constant as $n$ grows [^3] Note that, as expected, when r=n1r=n-1 the KS CI indicate a clear departure from normality. This occurs because we recover the complete uMMD, which under null is non-Gaussian.. These findings validate Theorem 2 for $k=2$, which predicts that for $r=O(\log^{q}(n))$ (with fixed $q$), deviations from normality are small and diminish as $n$ increases.

### v2

@@ Abstract
U-statistics are a fundamental class of estimators that generalize the sample mean and underpin much of nonparametric statistics. Although extensively studied in both statistics and probability, key challenges remain: their high computational cost—addressed partly through incomplete U-statistics—and their non-standard asymptotic behavior in the degenerate case, which typically requires resampling methods for hypothesis testing. This paper presents a novel perspective on U-statistics, grounded in hypergraph theory and combinatorial designs. Our approach bypasses the traditional Hoeffding decomposition, the main analytical tool in this literature but one highly sensitive to degeneracy. By characterizing the dependence structure of a U-statistic, we derive a Berry–Esseen bound valid for incomplete U-statistics of deterministic designs, yielding conditions under which Gaussian limiting distributions can be established even in degenerate cases and when the order diverges. We also introduce efficient algorithms to construct incomplete U-statistics of equireplicate designs, a subclass of deterministic designs that, in certain cases, achieve minimum variance. Finally, we apply our framework to kernel-based tests that use Maximum Mean Discrepancy (MMD) and Hilbert–Schmidt Independence Criterion. In a real data example with the CIFAR-10 dataset, our permutation-free MMD test delivers substantial computational gains while retaining power and type I error control.
@@ 1 Introduction
On the applications side, U-statistics have found widespread use across various fields. These range from classical statistical estimation and inference tasks: e.g., in nonparametric hypothesis testing (Bergsma and Dassios,, 2014; Yao et al.,, 2018), robust statistics (Joly and Lugosi,, 2016; Minsker and Wei,, 2020), bootstrap and resampling methods (Leucht and Neumann,, 2013; Bastian et al.,, 2024), to modern machine learning applications: e.g., in empirical risk minimization (Clémençon et al.,, 2008; Chen et al.,, 2023), supervised learning ensembles (Mentch and Hooker,, 2016; Peng et al.,, 2022) and kernel methods (Gretton et al.,, 2005, 2012; Liu et al.,, 2016). Second-order U-statistics are the most common and widely used in practice. This class includes estimators of variance, covariance, Kendall’s Tau, Spearman’s Rho, Wilcoxon statistics, Cramer-Von Mises statistics. Additionally, they encompass novel estimators of the Maximum Mean Discrepancy (MMD) and the Kernel Stein Discrepancy, which are kernel methods designed for two-sample testing (Gretton et al.,, 2012) and goodness-of-fit testing (Liu et al.,, 2016), respectively (see Schrab et al., (2022) for a discussion). Higher-order U-statistics are also relevant in applications, as they include estimators of dependence measures such as distance covariance (Székely et al.,, 2007), sign covariance (Bergsma and Dassios,, 2014), and the Hilbert–Schmidt Independence Criterion (HSIC) (Gretton et al.,, 2005).
@@
    We show that the dependency graph (Baldi and Rinott,, 1989) of an incomplete U-statistic is the line graph of the design hypergraph.
@@
Organization. Section 2 reviews background material and sets the notation for the paper. Section 3 derives a Berry–Esseen bound for incomplete U-statistics of deterministic designs and establishes conditions for Gaussian limiting distributions, even in the degenerate case and when the order diverges. Section 4 presents efficient algorithms for constructing equireplicate designs. Section 5 validates our theoretical results through extensive simulations on kernel-based test statistics and a real-data application showcasing a permutation-free variant of the MMD test. Section 6 concludes with a summary and directions for future research. All proofs and technical details corresponding to each section of the paper are provided in the supplementary material, where sections are labeled with an “S”. The R code for this paper can be found at https://github.com/CaesarXVII/IUstat_equireplicate_designs.
@@ 2.2 Equireplicate designs
When $k=2$, incomplete U-statistics based on equireplicate designs achieve minimum variance, as demonstrated by Theorem 1 on page 195 of Lee, (1990). When $k>2$, additional conditions are needed to ensure minimum variance, such as balanced incomplete block designs (BIBDs) and cyclic designs (see Lee, (1982), Examples 2 and 7, respectively). Further details are provided in S8.1.
The existence of equireplicate designs requires specific divisibility conditions. In general, given a nonzero integer $a$ and an integer $b$, we write $a\mid b$ to indicate that $a$ divides $b$, and $a\nmid b$ otherwise. For our design constructions, we use modular arithmetic and some basic group theory (see Gallian, (2021), ch. 2 for an overview). We write $b\pmod{n}$ to denote the remainder upon dividing $b$ by $n$, but set $b\pmod{n}\vcentcolon=n$ if the remainder is zero; we also write $\text{gcd}(b,n)$ as the greatest common divisor of $b$ and $n$. In Section 4, we present efficient algorithmic constructions of equireplicate designs that use these algebraic concepts. Since we relabeled 0 as $n$ when calculating values $\pmod{n}$, we consider $\mathbb{Z}_{n}=\{1,2,\ldots,n\}$ as the group of integers modulo $n$ with additive identity element $n$. This notation was chosen to ensure that the elements of our designs are labeled from the index set $V=\{1,\ldots,n\}$.
@@ 3.1 The dependency graph of incomplete U-statistics
In a dependency graph, there is an edge between two random variables only if they are dependent (absence of an edge implies independence). In the case of U-statistics, we include an edge between two random variables $h(S_{1})$ and $h(S_{2})$ if and only if $S_{1}\cap S_{2}\neq\emptyset$.
@@
The maximum degree of the dependency graph $\Delta(L(\mathcal{D}))$, bounds how many random variables each $h(S)$ can be dependent with. Controlling this quantity is key for establishing a normal limiting distribution, as we demonstrate in the following section.
@@ 3.3 Asymptotic results in the finite and infinite order regimes
Moreover, from the Berry–Esseen bound in (8) with $p=3$, it follows that under the assumptions of Theorem 2, the convergence rate to the standard normal distribution is $n^{-1/2}$, up to logarithmic factors (see the proof of Theorem 2 for details). This rate matches the classical Berry–Esseen bound for the complete U-statistic in the finite-order non-degenerate case (see Lee, (1990), ch. 3.3 and references therein), again up to logarithmic terms. It is also only slightly slower than the $o(n^{-1/2})$ rate established in Korolyuk and Borovskikh, (1989) for the complete U-statistic in the finite-order degenerate case, where the limiting distribution is non-Gaussian. In that setting, the variance scales as $n^{k/2}$ for the complete statistic, while for the incomplete one it scales as $|D_{n}|^{1/2}$ (see Corollary 2). This last result confirms—and extends to deterministic designs—the findings of Remark 3.2 in Chen and Kato, (2019), which were stated for random designs.
@@
To apply Theorem 2 in the infinite-order regime, we need a consistent estimator of the variance when $k$ diverges. Existing works have proposed variance estimation methods for incomplete U-statistics in the infinite-order regime, which can be applied (Wang and Lindsay,, 2014; Song et al.,, 2019; Xu et al.,, 2024). However, specific details for implementation of our framework in this regime is left for future work.
@@ 4 Efficient Construction of Equireplicate Designs
In this section, we highlight another major advantage of equireplicate designs: they can be constructed in linear time with respect to the design size.
For second-order incomplete U-statistics, equireplicate designs of any given size can be efficiently constructed and have minimum variance. Therefore, equireplicate designs should be preferred when $k=2$, the most common and widely used case in practice.
When $k>2$, constructing equireplicate designs becomes challenging, as it relates to open problems in discrete mathematics. In Section 4.2 we present a construction which has still a linear computational complexity in the design size—when $k$ is fixed—and that allows $|D|=O(n^{2})$ (see Remarks S8 and 7 for further details). This construction, based on cyclic permutations (see Lee, (1982), Example 7), may be of independent interest for the combinatorial design and hypergraph community.
4.1 Construction of r-equireplicate designs when k = 2
@@ 5.1 MMD experiments
We observe that for $r\in\{\log(n),\,\log^{2}(n)\}$, the KS CIs lie below $Q_{0.975}$ even at $n=100$ and quickly contract toward $Q_{0.5}$ as $n$ increases. When $r=\log^{3}(n)$, a larger sample size is required before the upper bound drops below $Q_{0.975}$, but the decreasing trend is evident. In contrast, for larger values of $r$ (e.g., $r=n/2$), the distance to normality remains roughly constant as $n$ grows [^3] Note that, as expected, when r=n1r=n-1 the KS CI indicate a clear departure from normality. This occurs because we recover the complete uMMD, which under the null is non-Gaussian.. These findings validate Theorem 2 for $k=2$, which predicts that for $r=O(\log^{q}(n))$ (with fixed $q$), deviations from normality are small and diminish as $n$ increases.